A ream of paper has 500 pages and the stack of paper is 4.5 × 10^(-2 metres high.
Find the thickness of 1 sheet of paper. State the number of significant figures in
your answer
1
Expert's answer
2020-08-24T19:07:59-0400

Given that


The height of stack/ream of paper


h = 4.5x10-2 meters

h = 4.5x10-2 x 103 mm

h = 45 mm (millimetres)


Total number of papers in the stack n = 500


So the thickness of each sheet t = h÷n = 45 ÷ 500

t = 0.09 mm